# Hybrid deep learning-based strategy for the hepatocellular carcinoma cancer grade classification of H&E stained liver histopathology images

### v1

@@ II-C Training
Furthermore, the cross entropy loss $(L_{CE})$ between labels and ground truth values is used, which is defined in Eq.[2]
$L_{CE}=-\sum_{i=0}^{N-1}y_{i}\log(p_{i})$ (2)
where $N$ is the number of classes, and $y_{i}$ and $p_{i}$ are the ground truth values and the predicted probabilistic values, respectively. This loss function is widely preferred over the others. The reason for this is that, if we see the curve of loss vs. probability, loss is very high when prediction is inaccurate. As a result, the gradient of loss is high, which aids in faster convergence. Once the model is trained, the training weights are used to transform the test dataset into the output cancer types using the test model. The performance of the model is evaluated on the test data using five performance measures such as accuracy, sensitivity, specificity, F1-score, and area under the curve (AUC). The pre-trained models were imported from the Pytorch library and trained on Google Colab.
Figure 5: The architecture of the proposed model.
@@
II-D Performance Metrics
The quantitative performance of the proposed hybrid model was evaluated using a standard protocol such as accuracy, sensitivity, specificity, and F1-score [39, 6]. These performance evaluation metrics were computed using a contingency table that holds positive (TP), false positive (FP), true negative (TN), and false negative (FN). Finally, receiver operating characteristics (ROC) and area under the curve (AUC) [40] were determined. li2019staged
III Results
@@
III-A Results on TCGA-LIHC dataset
@@ III-B Results on KMC dataset
Figure 10: Train and validation loss plotted for all epochs, all 5 folds over KMC train dataset of hybrid model with EfficientNetb3 as feature extractor
Figure 11: Train and validation accuracy plotted for all epochs, all 5 folds over KMC train dataset of hybrid model with EfficientNetb3 as feature extractor
@@
III-C Results on COLON dataset
@@ IV Discussion
The goal of this study is to develop a methodology for classifying different types of liver HCC tumors. We designed a deep learning-based hybrid architecture, trained and tested over two liver HCC datasets. The different data pre-processing steps along with training and validation techniques are also incorporated to build a robust model. After training the models, performance metrics over the test dataset were generated and observed. In this section, conclusions made from the obtained results, a comparison of the proposed methodology with recent studies and limitations of the proposed methodology are mentioned.
@@
We observe that hybrid models perform better than base models for all datasets mentioned above. Hybrid models show a minimum 0.76% (VGG16) and 3.29% (VGG16) of hike in accuracy on TCGA and KMC datasets respectively. The hybrid model with ResNet50 as a feature extractor outperforms the others on TCGA, achieving 100% accuracy with a 1.74% increase after modification (Tables III, IV) for TCGA dataset. Performance metrics indicate the excellent ability of classification and the robustness of the model. The hybrid model with EfficientNetb3 outperforms all other models on KMC in terms of every performance metric (Tables V, VI). With an accuracy of 96.71% and a 2.22% standard deviation, this model sets the benchmark for KMC database. A similar trend is observed for the COLON dataset as well. The hybrid models with EfficientNetb2, EfficientNetb4 and DenseNet121 feature extractors show remarkable performance by achieving 100% accuracy on the test dataset. These results prove that replacing a shallow classifier with a deep classifier and fine-tuning the top layers of the feature extractor is a fruitful strategy.
@@
For the TCGA dataset, the base model with EfficientNetb0 as a feature extractor gives better accuracy, whereas, among the hybrid models, ResNet50 is proven to be the best. On the other hand, for the KMC dataset, EfficientNetb3 outperforms other architectures. Various architectures are observed to produce the best performance across the datasets. This concludes the data specificity of deep architectures and not model specificity. Various architectures are designed in a way to identify or detect certain types of features that are associated with particular datasets, even when the tissues are those of the liver, but share different genetic patterns and different ways of WSI preparation procedures.
@@
Some recent studies on liver HCC detection involving the TCGA and KMC datasets are mentioned in Table IX. As can be seen, the proposed methodology outperforms other studies on local image classification tasks in terms of accuracy as well as other performance metrics. LiverNet architecture [2] claimed the best performance over KMC and TCGA datasets in the patch-level classification task. The LiverNet model has some key features such as CBAM, ASPP block and hypercolumn technique, which take care of a lesser number of model parameters and show improved performance. Our proposed methodology exceeds LiverNet’s accuracy by more than 2% over the TCGA dataset and by more than 5% on the KMC dataset, despite having less number of patches in the KMC dataset. LiverNet achieves F1-score of 97.72% and 90.93%. On the other hand, our methodology achieves 100% and 96.446% F1-score on TCGA and KMC respectively. They had also produced results using a very similar architecture called BreastNet[20] which was specifically designed for breast cancer classification over breakhis dataset[41]. Our proposed model outperforms BreastNet as well.
@@
First Author(year) Dataset Types Model Metrics Training method Remark Sun et al. TCGA 2 Pretrained Accuracy 100 10 fold Transfer learning, (2019)[12] 189,531 train, ResNet50 Precision 100 CVb k feature selection patches Recall 100 from sorted aggregated patch 224x224 F1-score 100 features, global level tested over 70WSI classification Chen et al. TCGA 3 Pretrained Accuracy 89.6 0.85:0.15 train Transfer learning, (2020)[13] 39k train, Inceptionv3 Precision 87.9 test split EasyDL framework patches, 9k test Recall 77.1 256 × 256 F1-score 82.0 very big train set MCCa 0.912 Huge number of paramenters Aatresh et al. TCGA 3 LiverNet Accuracy 97.72 5 fold Architecture specific (2021)[2] 2240 train Precision 97.72 CV for HCC liver cancer classification. 140 test Recall 97.72 CBAM, ASPP block 224x224 F1-score 97.72 hypercolumn technique IOUc 95.61 less number of parameters Aatresh et al. TCGA 3 BreastNet Accuracy 96.04 5 fold Architecture specific (2021)[2] 2240 train Precision 96.04 CV for Breast cancer classification. 140 test Recall 96.04 CBAM block hypercolumn 224x224 F1-score 96.04 technique IOU 92.47 less number of parameters Aatresh et al. KMC 2 LiverNet Accuracy 90.93 5 fold Architecture specific (2021)[2] 3210 train Precision 90.93 CV for HCC Liver cancer classification. 272 test Recall 90.93 CBAM, ASPP block 224x224 F1-score 90.93 hypercolumn technique IOU 83.6 Aatresh et al. KMC 2 BreastNet Accuracy 88.41 5 fold Architecture specific (2021) [2] 3210 train Precision 88.41 CV for Breast cancer classification. 272 test Recall 88.41 CBAM block hypercolumn 224x224 F1-score 88.41 technique IOU 79.73 Proposed TCGA 3 Pretrained Accuracy 100.00 5 fold Transfer learning, fine Hybrid 3528 train Hybrid Sensitivity 100.00 CV tuning, deep classifier ResNet50 392 test ResNet50 Specificity 100.00 training of selective (2022) 224x224 F1-score 100.00 top layer of base model AUC 1.00 Proposed KMC 2 Pretrained Accuracy 96.71 5 fold Transfer learning, fine hybrid 2725 train Hybrid Sensitivity 96.97 CV tuning, deep classifier EfficientNetb3 280 test EfficientNetb3 Specificity 98.85 training of selective (2022) 224x224 F1-score 96.71 top layer of base model AUC 0.99 aMCC stands for Matthews’s correlation coefficient. bCV stands for Cross Validation. bIOU stands for intersection over union
@@
For 3-class patch-level classification on TCGA and 2-class patch-level classification on KMC datasets, the proposed hybrid methodology outperforms all the above-mentioned studies. Furthermore, the proposed model achieves better performance despite being trained on a comparatively smaller dataset and provides end-to-end deep learning-based solutions. The performance metrics also indicate the robustness of proposed models over others. We have also demonstrated the performance of our methodology over colon histopathology images. Here as well hybrid models have achieved 100% accuracy.
There are some limitations to this study. ResNet50, EfficientNetb3, and DenseNet121-based architectures have around 25M, 12.5M and 7M parameters, respectively. These models have a large number of parameters, which makes them hard to run on low-end processor specifications. We will try to overcome this limitation in future work and provide a highly accurate and lightweight solution.
V Conclusion
@@
This study was focused on liver HCC classification using histopathology images. The primary dataset TCGA-LIHC was prepared through pre-processing steps involving patch extraction, color normalization and augmentation techniques. Another proprietary KMC dataset and a publically available COLON dataset were used for validation purposes. Through this study, we proposed a deep learning-based hybrid architecture for the classification of liver HCC efficiently and robustly. From the experimental results, it is also observed that the proposed hybrid model outperforms the base model and other state-of-the-art models by a significant margin. The best results on the TCGA dataset were obtained by a hybrid model with ResNet50 as the feature extractor, giving an accuracy of 100$\pm$0.00 whereas on the KMC dataset, the proposed hybrid model with EfficientNetb3 as a feature extractor achieves an accuracy of 96.71$\pm$0.68%. Moreover, the proposed hybrid models provide outstanding results on the COLON dataset as well. The results lead us to the conclusion that deepening the classifier and customizing the top layers of the feature extractor can help improve performance significantly.

### v2

@@ III-A Settings
Figure 5: The workflow of the training and validation of the proposed method
@@
III-B Performance Metrics
The quantitative performance of the proposed hybrid model was evaluated using a standard protocol such as accuracy, sensitivity, specificity, and F1-score [39, 6]. These performance evaluation metrics were computed using a contingency table that holds true positive (TP), false positive (FP), true negative (TN), and false negative (FN) rates. Finally, receiver operating characteristics (ROC) and area under the curve (AUC) [40] were determined.
III-C Results
@@
III-C1 Results on the TCGA-LIHC dataset
@@ III-C2 Results on KMC dataset
Figure 10: Training and validation loss plotted for all epochs, all 5 folds over KMC train dataset of hybrid model with EfficientNetb3 as feature extractor
Figure 11: Training and validation accuracy plotted for all epochs, all 5 folds over KMC train dataset of hybrid model with EfficientNetb3 as feature extractor
@@
III-C3 Results on COLON dataset
@@ IV Discussions
The goal of this study is to develop a methodology for classifying different types of liver HCC tumors. We proposed a deep learning-based hybrid architecture, trained and tested over two liver HCC datasets. The different data pre-processing steps along with training and validation techniques are also incorporated to build a robust model. After training the models, performance metrics over the untouched test dataset were generated and observed. In this section, conclusions made from the obtained results, a comparison of the proposed methodology with recent studies and limitations of the proposed methodology are mentioned.
@@
We observe that the hybrid model performs better than the base model for all datasets mentioned above. The hybrid model shows a minimum 0.76% (VGG16) and 3.29% (VGG16) of hike in accuracy on the TCGA and KMC datasets, respectively. The hybrid model with ResNet50 as a feature extractor outperforms the others on TCGA, achieving 100% accuracy with a 1.74% increase after modification (Tables III, IV) for the TCGA dataset. Performance metrics indicate the excellent ability of the classification and the robustness of the model. The hybrid model with EfficientNetb3 outperforms all other models on KMC in terms of every performance metric (Tables V, VI). With an accuracy of 96.71% and a 2.22% standard deviation, this model sets the benchmark for KMC database. A similar trend is observed for the COLON dataset as well. The hybrid models with EfficientNetb2, EfficientNetb4 and DenseNet121 feature extractors show remarkable performance by achieving 100% accuracy on the test dataset. These results prove that replacing a shallow classifier with a deep classifier and fine-tuning the top layers of the feature extractor is a fruitful strategy.
@@
For the TCGA dataset, the base model with the EfficientNetb0 as a feature extractor gives better accuracy, whereas, among the hybrid models, ResNet50 is proven to be the best. On the other hand, for the KMC dataset, EfficientNetb3 outperforms other architectures. Various architectures are observed to produce the best performance across the datasets. This concludes that, model performance is inherently dependent on dataset-specific characteristics, precluding the existence of a universally optimal architecture across diverse data distributions. Various architectures are designed in a way to identify or detect certain types of features that are associated with particular datasets, even when the tissues are those of the liver, but share different genetic patterns and different ways of WSI preparation procedures.
@@
Some recent studies on liver HCC detection involving the TCGA and KMC datasets are mentioned in Table IX. As can be seen, the proposed methodology outperforms other studies on local image classification tasks in terms of accuracy as well as other performance metrics. LiverNet architecture [2] claimed the best performance over KMC and TCGA datasets in the patch-level classification task. The LiverNet model has some key features such as CBAM, ASPP block and hypercolumn technique, which take care of a lesser number of model parameters and show improved performance. Our proposed methodology exceeds LiverNet’s accuracy by more than 2% over the TCGA dataset and by more than 5% on the KMC dataset, despite having less number of patches in the KMC dataset. LiverNet achieves F1-score of 97.72% and 90.93%. On the other hand, our methodology achieves 100% and 96.446% F1-score on the TCGA and KMC, respectively. They had also produced results using a very similar architecture called BreastNet[20], which was specifically designed for breast cancer classification over BreaKHis dataset[41]. Our proposed model outperforms BreastNet as well.
@@
First Author(year) Dataset Types Model Metrics Training method Remark Sun et al. TCGA 2 Pretrained Accuracy 100 10 fold Transfer learning, (2019)[12] 189,531 train, ResNet50 Precision 100 CVb k feature selection patches Recall 100 from sorted aggregated patch 224x224 F1-score 100 features, global level tested over 70WSI classification Chen et al. TCGA 3 Pretrained Accuracy 89.6 0.85:0.15 train Transfer learning, (2020)[13] 39k train, Inceptionv3 Precision 87.9 test split EasyDL framework patches, 9k test Recall 77.1 256 × 256 F1-score 82.0 very big train set MCCa 0.912 Huge number of paramenters Aatresh et al. TCGA 3 LiverNet Accuracy 97.72 5 fold Architecture specific (2021)[2] 2240 train Precision 97.72 CV for HCC liver cancer classification. 140 test Recall 97.72 CBAM, ASPP block 224x224 F1-score 97.72 hypercolumn technique IOUc 95.61 less number of parameters Aatresh et al. TCGA 3 BreastNet Accuracy 96.04 5 fold Architecture specific (2021)[2] 2240 train Precision 96.04 CV for Breast cancer classification. 140 test Recall 96.04 CBAM block hypercolumn 224x224 F1-score 96.04 technique IOU 92.47 less number of parameters Aatresh et al. KMC 4 LiverNet Accuracy 90.93 5 fold Architecture specific (2021)[2] 3210 train Precision 90.93 CV for HCC Liver cancer classification. 272 test Recall 90.93 CBAM, ASPP block 224x224 F1-score 90.93 hypercolumn technique IOU 83.6 Aatresh et al. KMC 4 BreastNet Accuracy 88.41 5 fold Architecture specific (2021) [2] 3210 train Precision 88.41 CV for Breast cancer classification. 272 test Recall 88.41 CBAM block hypercolumn 224x224 F1-score 88.41 technique IOU 79.73 Bhattacharya et al. LC25000 2 EfficientNeb4 Accuracy 99.99 5 fold AdBet-WOA, was applied (2023) [42] 5000 train F1-score 99.99 CV for feature selection, followed 500 test by classification using SVM 768x768 Talukder et al. LC25000 2 Ensemble Learning Accuracy 100.00 10 fold Deep extracted features fed (2022) [43] 5000 train F1-score 100.00 CV to 6 different algorithms, 500 test Precision 100.00 ensemble learning 768x768 Recall 100.00 AUC 1.00 Proposed TCGA 3 Pretrained Accuracy 100.00 5 fold Transfer learning, fine Hybrid 3528 train Hybrid Sensitivity 100.00 CV tuning, deep classifier ResNet50 392 test ResNet50 Specificity 100.00 training of selective 224x224 F1-score 100.00 top layer of base model AUC 1.00 Proposed KMC 4 Pretrained Accuracy 96.71 5 fold Transfer learning, fine Hybrid 2725 train Hybrid Sensitivity 96.97 CV tuning, deep classifier EfficientNetb3 280 test EfficientNetb3 Specificity 98.85 training of selective 224x224 F1-score 96.71 top layer of base model AUC 0.99 Proposed LC25000 2 Pretrained Accuracy 100.00 5 fold Transfer learning, fine Hybrid 5000 train Hybrid Sensitivity 100.00 CV tuning, deep classifier EfficientNetb2 500 test EfficientNetb2 Specificity 100.00 training of selective 768x768 F1-score 100.00 top layer of base model AUC 1.00 aMCC stands for Matthews’s correlation coefficient. bCV stands for Cross Validation. bIOU stands for intersection over union
@@
For 3-class patch-level classification on TCGA and 4-class patch-level classification on KMC datasets, the proposed hybrid methodology outperforms all the above-mentioned studies. Furthermore, the proposed model achieves better performance despite being trained on a comparatively smaller dataset and provides end-to-end deep learning-based solutions. The performance metrics also indicate the robustness of the proposed model over others. We have also demonstrated the performance of our methodology over colon histopathology images. Previously, Bhattacharya et al. [42] and Talukder et al. [43] on the colon dataset claimed very large accuracies at the cost of large computational expenses. Whereas Hybrid models have achieved 100% accuracy despite having a comparatively smaller number of parameters and end-to-end solution. Here as well, Hybrid models achieved 100% accuracy.
There are some limitations to this study. The architectures based on ResNet50, EfficientNetb3, and DenseNet121 have approximately 25M, 12.5M, and 7M parameters, respectively. The high parameter count of these models makes them challenging to operate on low-end processors. To address this issue, we plan to utilize distillation methods or explore alternative approaches using low-rank deep learning models in future work, aiming to deliver highly accurate and lightweight solutions.
V Conclusions
@@
This study was focused on liver HCC classification using histopathology images. The primary dataset TCGA-LIHC was prepared through pre-processing steps involving patch extraction, color normalization and augmentation techniques. Another proprietary KMC dataset and a publically available COLON dataset were used for validation purposes. Through this study, we proposed a deep learning-based hybrid architecture for the efficient and robust classification of liver HCC. The experimental results also show that the proposed hybrid model outperforms the base model and other state-of-the-art models by a significant margin. The best results on the TCGA dataset were obtained by a hybrid model with ResNet50 as the feature extractor, giving an accuracy of 100$\pm$0.00 whereas on the KMC dataset, the proposed hybrid model with EfficientNetb3 as a feature extractor achieves an accuracy of 96.71$\pm$0.68%. Moreover, the proposed hybrid models provide outstanding results on the COLON dataset as well. The findings suggest that enhancing the classifier, optimizing its depth, and selectively fine-tuning the top layers of the feature extractor can significantly boost performance.
@@
Acknowledgements
This work was funded by the Science and Engineering Research Board, India under Grant SERB- CRG/2021/005752.